OpenStudy (anonymous):

-5+14=6b-8-5b= help solve. thanks. :-)

Directrix (directrix):

Is this the problem: -5+14=6b-8-5b ? Is the task to solve for the variable b?

OpenStudy (anonymous):

solve

Directrix (directrix):

-5+14=6b-8-5b --------------- This is a linear equation. Simplify each side. That is, -5+14= ? and then 6b - 8 - 5b = ??

OpenStudy (anonymous):

As directrix said, linear. Solve for each side then solve for b. So -5+14=9 6b-8-5b=1b-8 So problem becomes 9=1b-8 Then you solve for b by adding 8 to both sides: 9+8=17 on one side 1b-8+8 =1b on the other side Now it's 17=1b so you divide by 1 on each side. 17/1=1b/1 So answer as requested is b=17

OpenStudy (anonymous):

-5+14=6b-8-5b Since b is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 6b-8-5b=-5+14 Since 6b and -5b are like terms, add -5b to 6b to get b. b-8=-5+14 Add 14 to -5 to get 9. b-8=9 Since -8 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 8 to both sides. b=8+9 Add 9 to 8 to get 17. b=17